Calculate the relative rates of effusion of He(g) and O2(g).


1
Expert's answer
2021-11-06T01:29:57-0400

If the two gases A and B are at same temperature and pressure, the ratio of their diffusion rates is equal to the inverse of ratio of the square roots of the masses of their particles or molecular masses.

"\frac{rate \;of \;diffusion\; of\; gas\; A}{rate\; of\; diffusion\; of\; gas\; B} = \frac{\sqrt{M_B}}{\sqrt{M_A}}"

M(He) = 4.0 g/mol

M(O2) = 32.0 g/mol

"\frac{rate \;of \;diffusion\; of \; He}{rate\; of\; diffusion\; of\; O_2} = \frac{\sqrt{32.0}}{\sqrt{4.0}} \\

= \frac{5.656}{2} = 2.828"